### v1

# A Two-Step Minimax Q-learning Algorithm for Two-Player Zero-Sum Markov Games

Abstract
An interesting iterative procedure is proposed to solve a two-player zero-sum Markov games. First this problem is expressed as a min-max Markov game. Next, a two step Q-learning algorithm for solving Markov decision problem (MDP) is suitably modified to solve this Markov game. Under a suitable assumption, the boundedness of the proposed iterates is obtained theoretically. Using results from stochastic approximation, the almost sure convergence of the proposed two-step minimax Q-learning is obtained theoretically. More specifically, the proposed algorithm converges to the game theoretic optimal value with probability one, when the model information is not known. Numerical simulation authenticate that the proposed algorithm is effective and easy to implement.
Index Terms:
Two-player zero-sum Markov games, Multi-agent reinforcement learning, minimax Q-learning.
11footnotetext: Department of Mathematics, Indian Institute of Technology (IIT) Indore, India, 452020. Email:shreyassr123@gmail.com, phd1901241006@iiti.ac.in, Shreyas S R is sincerely grateful to CSIR-HRDG, India (Grant No. 09/1022(0088)2019-EMR-I) for the financial support22footnotetext: Department of Mathematics, IIT Indore, Simrol, Indore-452020, Madhya Pradesh, India. Email:vijesh@iiti.ac.in
Single-agent reinforcement learning (RL) deals with the task of finding an optimal policy or strategy in a stochastic environment. Markov Decision Process (MDPs) provides the underlying framework for solving the problem of finding an optimal strategy in these decision making problems with one agent [1]. Research in RL showcases the popularity of single-agent RL algorithms. Particularly, single-agent RL has been successfully implemented in robot navigation [2], agent based production scheduling [3], economics [4], and traffic signal control [5]. However, many practical environments involve more than one agents competing against each other to optimize their strategies. For example in real life, consumers and sellers are competing to find optimal price. This type of problem can be studied systematically using the concept of Markov games from game theory. Markov games can be considered as a generalisation of MDPs [6]. Similar to the problem of finding an optimal strategy in MDPs, the problem of finding an optimal strategy in Markov games is an important problem. Using successive iterative scheme, when the model information is completely available, Shapley [7] obtained the optimal strategy for two-player zero-sum Markov games (TZMG). More specifically, Shapley [7] solved a non-linear equation involving minimum and maximum operator. This non-linear equation is also called as the Shapley equation [8, 9]. This equation is similar to the Bellman’s equation arising from MDPs. The Q-learning algorithm from stochastic approximation enable the users to find an optimal strategy in Markov decision problem even if the model information is not completely available. Using this, M. L. Littman [10] extend the Q-learning algorithm to solve TZMG when the information of the model is not completely available. This algorithm is also called as minimax Q-learning. Under suitable assumptions M. L. Littman and C. Szepesvari [11] show that this algorithm converges to the game theoretic optimal value with probability one. Consequently, Markov games become an elegant framework for reinforcement learning with more than one agent, though efforts were made without Markov games [12, 13].
The Q-learning algorithm for TZMG [10] is further designed to handle both coordination games [14], and general sum games [15]. The Q-learning algorithm for general sum stochastic game [15] was refined by M. Bowling [16] and further generalised by J. Hu and M. P. Wellman [17]. This algorithm in [17] is called as Nash Q-learning. More specifically, the Q-values from the Nash-Q-learning algorithm converges to Nash Q-values with probability one under some restrictive conditions. By introducing the attribution friend and foe, M. L. Littman [18] relaxed the restrictive assumption in the Q-learning algorithm for general sum stochastic games [15]. By introducing WoLF principle, M. Bowling and M. Veloso [19, 20] tried to develop multi-agent reinforcement learning satisfying rational property, using the stochastic game framework. Minimax Q-learning algorithm together with value function approximation was studied by M. G. Lagoudakis and R. Parr [21] for TZMG. Further, A. Greenwald and K. Hall [22] developed an algorithm called correlated Q-learning for solving general sum Markov games. This algorithm can be considered as a generalisation of Littman’s [18] friend-or-foe Q-learning and Hu and Wellman’s [17] Nash Q-learning. In this direction, various other reinforcement algorithms also suitably modified to study TZMG. B. Banerjee et al. [23] extended the SARSA algorithm as minimax-SARSA algorithm to find Nash equilibrium in TZMG. J. Perolat et al. [24] customised various non-stationary reinforcement learning algorithms to the discounted zero-sum Markov games. A stochastic approximation algorithm based on fictitious play and actor-critic algorithm was proposed by J. Perolat et al. [25] to find Nash equilibrium in zero-sum two-player multi-stage game as well as cooperative multi-stage games. Efforts were made in [26], and [27] to study the sample complexity in TZMG algorithms. Error propagation in algorithms from approximate dynamic programming for TZMG were studied by Perolat et al. [28]. Interesting algorithms based on minimax Q-learning together with deep Q-network for studying TZMG is available in recent literature [29, 30]. It is worth to mention that a theoretical analysis is provided in [29] for minimax deep Q-network algorithm for TZMG. The SOR Q-learning [31] for MDP is suitably updated to handle TZMG in [32]. Similarly, the switching system technique for MDP [33] was modified to study minimax Q-learning for TZMG [34]. Various developments in the aspects of multi-agent reinforcement learning algorithms have been comprehensively reported in [35][36].
Though various single-step RL algorithms with single-agent suitably updated for multi-agent RL algorithms in literature, very few results are available for multi-step multi-agent RL algorithms. In 2001, B. Banerjee et al. [23] observed that Peng’s and William’s Q($\lambda$) [37] version of minimax algorithm performs better than the single-step SARSA minimax [23] and Littman’s minimax [10] algorithms. Unfortunately, they were not able to obtain the theoretical convergence of this multi-step multi-agent RL algorithm. It is worth to mention that the theoretical convergence of Peng’s and William’s Q($\lambda$) algorithm for single-agent was also not available during this period. Recently, in 2021, Kozuno et al. [38], proved the convergence of Peng’s and William’s Q($\lambda$) algorithm for a single-agent. Studying convergence of multi-step algorithms for MDP are available in [38, 39, 40]. In [41], two interesting two-step algorithms were proposed to handle MDP and obtained their theoretical convergence results. Numerical results in [41] showed that the two-step method has good flexibility and performs better than the single-step methods in some benchmark problems. In this manuscript we customize this idea from [41] to handle Markov games. More specifically, a two-step minimax Q-learning (TMQL) algorithm is proposed to obtain the Nash equilibrium in the TZMG. The contributions of this manuscript are as follows :
  • The boundedness of the proposed TMQL algorithm is provided.
  • Using the techniques available in the stochastic approximation theory the almost sure convergence of the TMQL algorithm is presented.
  • The proposed algorithms performance is empirically compared with recent algorithms, demonstrating the superiority of the proposed algorithm both in terms of error and time.
0-A Related works:
Earlier literature explored the notion of multi-step return in the context of prediction problem [42]. Further, Watkins delves into the concept of multi-step returns and the trade-off between bias and variance within the single-agent framework [43]. Several multi-step algorithms [37, 43, 44] were proposed towards the end of twentieth century without rigorous convergence analysis. Lately, many of the open questions regarding the convergence of previously proposed multi-step algorithms are discussed in [39], and [40]. More specifically, the convergence of multi-step algorithms for control problems like Watkin’s Q($\lambda$), Peng’s and William’s Q($\lambda$), and Naive Q($\lambda$) are discussed in [40], [38], and [39] respectively. Further, new multi-step algorithms are also getting added in this list, for example the Retrace($\lambda$) [40], and Q($\sigma$) [45] algorithms. The practicality of employing multi-step algorithms have long been a subject of inquiry. This question finds a comprehensive response in [46], and [47]. In other words, [46], and [47] explains the advantages offered by multi-step algorithms by conducting extensive empirical tests. While numerous multi-step algorithms have been developed within the framework of single-agent RL, there is still much work to be done in exploring their application and advantages within the context of MARL. However, recent success of algorithms like AlphaGo which uses multi-step Monte Carlo Tree Search (MCTS) algorithm is an important step in that direction [48]. In a similar vein the proposed algorithm also explores the notion of multi-step RL algorithms in the context of TZMG.
The structure of the paper is as follows: To make the manuscript self-contained, Section 2 covers preliminaries, notations, and essential results. Section 3 discusses the proposed algorithm in detail. Section 4 presents the main findings and demonstrates the almost sure convergence of the proposed algorithm. Section 5 includes numerical experiments. Finally, Section 6 draws conclusions.
1 Preliminaries and Background
In this manuscript, we consider a two-player zero-sum Markov games. Formally, described by a six-tuple $(S,A,B,p,r,\alpha)$, where $S$ is a set of states, $A$ and $B$ are the set of actions that can be performed by agent 1, and agent 2 respectively. The system evolves according to the transition probability $p$. More specifically, $p(j|i,a,b)$ represents the probability of transitioning to state $j$ when agent $1$, and agent $2$ select actions $a$, and $b$, respectively, while in state $i$. The real number $r(i,a,b)$ denotes the reward or payoff obtained by player 1 when the actions $a$, and $b$ are taken by agent 1, and agent 2 respectively, while in state $i$. In zero-sum Markov games, it is important to note that the payoff for agent 2 is negative of the payoff acquired by agent 1. Further, $\alpha\in[0,1)$ denotes the discount factor. It is important to mention that this manuscript deals with finite state space $S$, and the finite action spaces $A$, and $B$. Also, the norm used is the max-norm, i.e., for any $z\in\mathbb{R}^{d}$, $\|z\|:=\max_{1\leq i\leq d}|z(i)|$. In addition, let $R_{\max}$ denote $\max_{i,a,b}|r(i,a,b)|$.
The objective of the agents in TZMG is to learn an optimal strategies $\mu^{*}_{1}:S\rightarrow\Delta^{|A|}$, and $\mu^{*}_{2}:S\rightarrow\Delta^{|B|}$, where $\Delta^{|X|}$ denote the probability simplex in $\mathbb{R}^{|X|}$. The existence of such policies for the discounted finite state action Markov game is guaranteed in the classical game theory [7]. The problem of finding an optimal strategy $\mu^{*}_{1}$, and $\mu^{*}_{2}$ reduces to solving the following system of non-linear equations analogous to Bellman’s optimality equations from dynamic programming given by
$Q^{*}(i,a,b)=$ $r(i,a,b)+\alpha\sum_{j=1}^{|S|}p(j|i,a,b)\,val[Q^{*}(j)],$ (1)
$\quad\quad\hskip 56.9055pt\forall(i,a,b)\in S\times A\times B,$
where the $val$ operator is defined for any $|A|\times|B|$ matrix $M$ as follows:
$val[M]:=\min_{y\in\Delta^{|B|}}\max_{x\in\Delta^{|A|}}x^{t}My,\quad\text{for any}\,M.$ (2)
Similar to the Q-Bellman operator in MDPs, the minimax Q-Bellman operator follows
$H:\mathbb{R}^{|S\times A\times B|}\rightarrow\mathbb{R}^{|S\times A\times B|}$ is defined as
$(HQ)(i,a,b)=r(i,a,b)+\alpha\sum_{j=1}^{|S|}p(j|i,a,b)val[Q(j)].$ (3)
The subsequent analysis relies on the properties of the $val$ operator, which are presented in the following lemma.
Lemma 1.1 (Proposition G.4, in [6]).
Let $M=[m_{ij}]$, and $N=[n_{ij}]$ be two $|A|\times|B|$ matrices. Then the following holds true :
(i) $|val[M]-val[N]|\,\leq\max_{i,j}|m_{ij}-n_{ij}|=\|M-N\|.$
(ii) $|val[M]|\,\leq\|M\|.$
Using Lemma 1.1, one can conclude that, $H$ is a contraction mapping with respect to max-norm, and $\alpha$ is the contraction factor. Consequently, $H$ has a unique fixed point, say $Q^{*}$. From the optimal $Q^{*}$, an optimal policies $\mu_{1}^{*}$, and $\mu_{2}^{*}$ obtained as follows
$(\mu_{1}^{*}(i),\mu_{2}^{*}(i))\in\arg val[Q^{*}(i)],\quad\forall i\in S.$
M. L. Littman and C. Szepesvari [11] proved that the following iterative scheme:
$Q_{n+1}(i,a,b)=(1-\beta_{n}(i,a,b))Q_{n}(i,a,b)$
$\quad\quad\quad+\beta_{n}(i,a,b)(r(i,a,b)+\alpha\,val[Q_{n}(j)]),$
converges to the optimal Q-value if all combinations of states and actions are chosen infinitely often and the learning rate satisfies Robbins and Monro condition [49]. In other words, minimax Q-learning converges to the fixed point of minimax Q-Bellman operator.
The following lemma from stochastic approximation due to S. Singh et al. [50] is used to prove the almost convergence of the proposed two-step minimax Q-learning algorithm.
Lemma 1.2.
(Lemma 1, [50]) Let $X$ be a finite set. Let $\left(\Psi_{n},F_{n},\beta_{n}\right)$ be a stochastic process such that $\Psi_{n},F_{n},\beta_{n}:X\rightarrow\mathbb{R}$ and satisfies the recurrence relation :
[…]
1. $0\leq\beta_{n}(y)\leq 1$, $\sum_{n=1}^{\infty}\beta_{n}(y)=\infty$, $\sum_{n=1}^{\infty}\beta^{2}_{n}(y)<\infty$ w.p.1.
[…]
3. $Var[F_{n}(y)|\mathcal{F}_{n}]\leq K(1+\|\Psi_{n}\|)^{2}$, where $K$ is some constant.
Boundedness plays an important role in the convergence of many iterative schemes in RL. Usually by assuming the boundedness of the Q-iterates [32] or by proving the boundedness of the Q-iterates [50][51] the theoretical convergence analysis will be performed. In this manuscript, the boundedness of the Q-iterates proved theoretically under suitable assumption on rewards using the following lemma.
Lemma 1.3.
(Proposition 3.1, [52]) If $\sum_{n=1}^{\infty}|a_{n}|<\infty$, then the product $\prod_{n=1}^{\infty}(1+a_{n})$ converges.
With the above results at hand we present the proposed algorithm in the next section and its convergence behaviour in Section 4.
2 Proposed Algorithm
The proposed algorithm is presented in this section. As discussed in the previous section, for standard minimax Q-learning, at every step of the iteration, we receive a reward $r(i,a,b)$ and the next state $j$ at state $i$. In the proposed algorithm, however, at every step of the iteration, in addition to $r(i,a,b)$, and $j$, we receive $r(j,c,d)$, and a new state $k$ by taking an action $c$, and $d$ at the state $j$. Together with the learning rate $\beta_{n}$, we use a sequence $\theta_{n}$ such that $0\leq\theta_{n}\leq 1$, ${|\theta_{n}|}$rightarrow 0 monotonically, and $\sum_{n=1}^{\infty}\beta_{n}|\theta_{n}|<\infty$.
Algorithm 1 Two-step minimax Q-learning (TMQL)
0:  Initial Q-Vector: $Q_{0}$; Discount factor: $\alpha$, step size rule $\beta_{n}$, and sequence $\theta_{n}$, Number of iterations: $T$. A trajectory $\{(s_{n},a_{n},b_{n},r^{\prime}_{n},s^{\prime}_{n},a^{\prime}_{n},b^{\prime}_{n},r^{\prime}_{n+1},s_{n+1},a_{n+1},b_{n+1})\}^{\infty}_{n=0}$ where every $S\times A\times B$ triplet appears infinitely often.
1:  for $n=0,1,\cdots,T-1$ do
2:   For actions $a_{n}$, and $b_{n}$ at $s_{n}$
3:   Observe $s^{\prime}_{n}$, and $r^{\prime}_{n}=r(s_{n},a_{n},b_{n})$
4:   For $a^{\prime}_{n}$, $b^{\prime}_{n}$ at $s^{\prime}_{n}$
5:   Observe $s_{n+1}$, and $r_{n+1}=r(s^{\prime}_{n},a^{\prime}_{n},b^{\prime}_{n})$
6:   Update rule: $Q_{n+1}(s_{n},a_{n},b_{n})=(1-\beta_{n}(s_{n},a_{n},b_{n}))Q_{n}(s_{n},a_{n},b_{n})+\beta_{n}(s_{n},a_{n},b_{n})\Big{(}r^{\prime}_{n}+\alpha\,val[Q_{n}(s^{\prime}_{n})]+\alpha\,\theta_{n}(r_{n+1}+\alpha\,val[Q_{n}(s_{n+1})])\Big{)}$
7:  end for
8:  return  $Q_{T-1}.$
As it can be observed from the Algorithm 1, for samples $\{i,a,b,r(i,a,b),j,c,d,r(j,c,d),k\}$ generated at time step $n$, the update rule for TMQL is of the following form:
$Q_{n+1}(i,a,b)=(1-\beta_{n}(i,a,b))Q_{n}(i,a,b)$ (5)
$\hskip 28.45274pt+\beta_{n}(i,a,b)(r(i,a,b)+\alpha val[Q_{n}(j)]$
$\hskip 56.9055pt+\alpha\theta_{n}(r(j,c,d)+\alpha val[Q_{n}(k)])).$
3 Main Results and Convergence Analysis
The following section discusses the almost sure convergence of the proposed algorithm using Lemma 1.2, discussed in Section 2. Before presenting the main convergence theorem the boundedness of the two-step minimax Q-learning will be discussed using the approach used in two-step Q-learning for MDP [41]. For the sake of making this manuscript self-contained the boundedness property of the Q-iterates is presented here.
Lemma 3.1.
Let $Q_{n}(i,a,b)$ be the value corresponding to a state $i$ and action pair $(a,b)$ at $n^{th}$ iteration of TMQL. If $\|Q_{0}\|\leq\dfrac{R_{\max}}{1-\alpha}$, then $\|Q_{n}\|\leq M,\,\forall n\in\mathbb{N},\quad\text{where}\;M=\dfrac{R_{\max}}{1-\alpha}(1+\alpha|\theta_{0}|)\prod_{i=1}^{\infty}(1+\beta_{i}|\theta_{i}|\alpha^{2})$.
Proof.
We obtain this result using induction principle. Let $(i,a,b)\in S\times A\times B$, {dmath*} —Q_1(i,a,b)—=—(1-β_0(i,a,b))Q_0(i,a,b)+β_0(i,a,b)(r(i,a,b)+αval[Q_0(j)]
+αθ_0 (r(j,c,d)+αval[Q_0(k)] ) ) —
≤(1-β_0)∥Q_0∥ + β_0 (R_max + α∥Q_0∥ )+ β_0 αθ_0— (R_max+α∥Q_0∥)
≤Rmax1-α(1+αθ_0—). Hence, we have $\max_{(i,a,b)}|Q_{1}(i,a,b)|=\|Q_{1}\|\leq\dfrac{R_{\max}}{1-\alpha}(1+\alpha|\theta_{0}|)\leq M$. For $k=1,2,...,n$, assume that $\|Q_{k}\|\leq\dfrac{R_{\max}}{1-\alpha}(1+\alpha|\theta_{0}|)\prod_{i=1}^{k-1}(1+\beta_{i}|\theta_{i}|\alpha^{2})\leq M$ is true . The proof will be completed if $\|Q_{n+1}\|\leq\dfrac{R_{\max}}{1-\alpha}(1+\alpha|\theta_{0}|)\prod_{i=1}^{n}(1+\beta_{i}|\theta_{i}|\alpha^{2})$. Define $L=\dfrac{R_{\max}}{1-\alpha}(1+\alpha|\theta_{0}|)\prod_{i=1}^{n-1}(1+\beta_{i}|\theta_{i}|\alpha^{2})$. Now, {dmath*} —Q_n+1(i,a,b)—=—(1-β_n(i,a,b))Q_n(i,a,b)+β_n(i,a,b)(r(i,a,b)+αval[Q_n(j)]
+αθ_n (r(j,c,d)+αval[Q_n(k)] ) ) —
≤L (1+β_nα^2—θ_n—)-β_n L+β_n R_max + αβ_n L + β_n αθ_n—R_max. Now we substantiate that, $Z_{n}=-\beta_{n}L+\beta_{n}R_{\max}+\alpha\beta_{n}L+\beta_{n}\alpha|\theta_{n}|R_{\max}\leq 0.$ For
$Z_{n}=\beta_{n}\left((\alpha-1)L+R_{\max}+\alpha|\theta_{n}|R_{\max}\right)$
$=\beta_{n}\bigg{(}(\alpha-1)\dfrac{R_{\max}}{1-\alpha}(1+\alpha|\theta_{0}|)\prod_{i=1}^{n-1}(1+\beta_{i}|\theta_{i}|\alpha^{2})$
$\hskip 28.45274pt+R_{\max}+\alpha|\theta_{n}|R_{\max}\bigg{)}$
$\leq\beta_{n}\left(-R_{\max}(1+\alpha|\theta_{0}|)+R_{\max}+\alpha|\theta_{n}|R_{\max}\right)\leq 0.$
Therefore, $\left|Q_{n+1}(i,a,b)\right|\leq L(1+\beta_{n}\alpha^{2}|\theta_{n}|)$. Thus, $\max_{(i,a,b)}|Q_{n+1}(i,a,b)|=\|Q_{n+1}\|\leq\dfrac{R_{\max}}{1-\alpha}(1+\alpha|\theta_{0}|)\prod_{i=1}^{n}(1+\beta_{i}|\theta_{i}|\alpha^{2})\leq M$. Hence the result follows. ∎
Throughout this manuscript, we assume that the following property holds true.
Assumption: The Markov chain generated by all control policies is ergodic. Furthermore, for any policy, there exists a positive probability of selecting any action in any given state.
Theorem 3.2.
Given a finite Markov game as defined in Section 2, the TMQL algorithm given by the update rule:
$Q_{n+1}(i,a,b)=(1-\beta_{n}(i,a,b))Q_{n}(i,a,b)$ (6)
$\hskip 56.9055pt+\beta_{n}(i,a,b)(r(i,a,b)+\alpha val[Q_{n}(j)]$
$\hskip 85.35826pt+\alpha\theta_{n}(r(j,c,d)+\alpha val[Q_{n}(k)])).$
[…]
Proof.
The correspondence to Lemma 1.2 follows from associating $X$ with $S\times A\times B$, and define $\Psi_{n}(s,a,b)=Q_{n}(s,a,b)-Q^{*}(s,a,b)$. Let
$s_{0},a_{0},b_{0},r^{\prime}_{0},s^{\prime}_{0},a^{\prime}_{0},b^{\prime}_{0},r_{1},s_{1},a_{1},b_{1},r^{\prime}_{1},s^{\prime}_{1},a^{\prime}_{1},b^{\prime}_{1},r_{2}\cdots$
be the trajectory produced by the behaviour strategies. Using the above notation, the update rule can be written as
$Q_{n+1}(s_{n},a_{n},b_{n})=(1-\beta_{n}(s_{n},a_{n},b_{n}))Q_{n}(s_{n},a_{n},b_{n})$
$\hskip 85.35826pt+\beta_{n}(s_{n},a_{n},b_{n})(r^{\prime}_{n}+\alpha\,val[Q_{n}(s^{\prime}_{n})]$
$\hskip 113.81102pt+\alpha\,\theta_{n}(r_{n+1}+\alpha\,val[Q_{n}(s_{n+1})])),$
where $r^{\prime}_{n}=r(s_{n},a_{n},b_{n})$, and $r_{n+1}=r(s^{\prime}_{n},a^{\prime}_{n},b^{\prime}_{n})$. Thus, {dmath*} Ψ_n+1(s_n,a_n,b_n)=(1-β_n(s_n,a_n,b_n))Ψ_n(s_n,a_n,b_n)+β_n(s_n,a_n,b_n)(r’_n+α  val[Q_n(s’_n)]+
                          αθ_n (r_n+1+α val[Q_n(s_n+1)] )-Q^*(s_n,a_n,b_n)). Let
$F_{n}(s_{n},a_{n},b_{n})$
$=r(s_{n},a_{n},b_{n})+\alpha\,val[Q_{n}(s^{\prime}_{n})]+\alpha\,\theta_{n}(r(s^{\prime}_{n},a^{\prime}_{n},b^{\prime}_{n})$
$\hskip 56.9055pt+\alpha\,val[Q_{n}(s_{n+1})])-Q^{*}(s_{n},a_{n},b_{n}).$
As the two-step minimax Q-learning requires extra sample during the update, for our analysis, we define the sigma fields as follows. For $n=0$, let $\mathcal{F}_{0}=\sigma(\{Q_{0},s_{0},a_{0},b_{0},\beta_{0}\})$ and for $n\geq 1$, $\mathcal{F}_{n}=\sigma(\{Q_{0},s_{0},a_{0},b_{0},\beta_{0},r^{\prime}_{j-1},s^{\prime}_{j-1},a^{\prime}_{j-1},b^{\prime}_{j-1},r_{j},\beta_{j},s_{j},a_{j},b_{j}:1\leq j\leq n\})$. With this choice of $\mathcal{F}_{n}$, $\beta_{0}$, and $\Psi_{0}$ are $\mathcal{F}_{0}$ measurable, and $\beta_{n}$, $\Psi_{n}$, and $F_{n-1}$ are $\mathcal{F}_{n}$ measurable. Now we have,
$\left|E[F_{n}(s_{n},a_{n},b_{n})|\mathcal{F}_{n}]\right|$
$=|E[r(s_{n},a_{n},b_{n})+\alpha\,val[Q_{n}(s^{\prime}_{n})]+\alpha\,\theta_{n}(r(s^{\prime}_{n},a^{\prime}_{n},b^{\prime}_{n})$
$\hskip 85.35826pt+\alpha\,val[Q_{n}(s_{n+1})]-Q^{*}(s_{n},a_{n},b_{n})|\mathcal{F}_{n}]|$
$=\left|E\left[r(s_{n},a_{n},b_{n})+\alpha\,val[Q_{n}(s^{\prime}_{n})]-Q^{*}(s_{n},a_{n},b_{n})|\mathcal{F}_{n}\right]\right.$
$\hskip 56.9055pt\left.+\alpha\theta_{n}E\left[r(s^{\prime}_{n},a^{\prime}_{n},b^{\prime}_{n})+\alpha\,val[Q_{n}(s_{n+1})]|\mathcal{F}_{n}\right]\right|$
$\leq\left|E\left[r(s_{n},a_{n},b_{n})+\alpha\,val[Q_{n}(s^{\prime}_{n})]-Q^{*}(s_{n},a_{n},b_{n})|\mathcal{F}_{n}\right]\right|$
$\hskip 56.9055pt+\alpha|\theta_{n}|\left|E\left[\left(r(s^{\prime}_{n},a^{\prime}_{n},b^{\prime}_{n})+\alpha\,val[Q_{n}(s_{n+1})]\right)|\mathcal{F}_{n}\right]\right|$
$=|\sum_{j=1}^{|S|}p(j|s_{n},a_{n},b_{n})(r(s_{n},a_{n},b_{n})+\alpha val[Q_{n}(j)]$
$\hskip 99.58464pt-Q^{*}(s_{n},a_{n},b_{n}))|$
$\hskip 56.9055pt+\alpha|\theta_{n}|\left|E\left[\left(r(s^{\prime}_{n},a^{\prime}_{n},b^{\prime}_{n})+\alpha\,val[Q_{n}(s_{n+1})]\right)|\mathcal{F}_{n}\right]\right|$
$\leq\left|HQ_{n}(s_{n},a_{n},b_{n})-HQ^{*}(s_{n},a_{n},b_{n})\right|+\alpha\,|\theta_{n}|\left(R_{\max}+\alpha\|Q_{n}\|\right)$
$\leq\alpha\,\|\Psi_{n}\|+\zeta_{n},$
where $\zeta_{n}=\alpha\,|\theta_{n}|(R_{\max}+\alpha\|Q_{n}\|)$. As the Q-iterates are bounded, and $|\theta_{n}|$rightarrow 0, one can conclude that, $\zeta_{n}$rightarrow 0 as $n\rightarrow\infty$. Therefore, condition $(2)$ of Lemma 1.2 holds.
Now consider,
$Var[F_{n}(s_{n},a_{n},b_{n})|\mathcal{F}_{n}]$
$=E\left[\left(F_{n}(s_{n},a_{n},b_{n})-E[F_{n}(s_{n},a_{n},b_{n})|\mathcal{F}_{n}]\right)^{2}|\mathcal{F}_{n}\right]$
$\leq E[(r(s_{n},a_{n},b_{n})+\alpha val[Q_{n}(s^{\prime}_{n})]$
$\hskip 28.45274pt+\alpha\theta_{n}(r(s^{\prime}_{n},a^{\prime}_{n},b^{\prime}_{n})+\alpha val[Q_{n}(s_{n+1})]))^{2}|\mathcal{F}_{n}]$
$\leq\left(R_{\max}+\alpha\|Q_{n}\|+\alpha|\theta_{n}|\left(R_{\max}+\alpha\|Q_{n}\|\right)\right)^{2}$
$\leq 3\left(R_{\max}^{2}+\alpha^{2}\|Q_{n}\|^{2}+\alpha^{2}(R_{\max}+\alpha\|Q_{n}\|)^{2}\right)\;$
$\leq 3R_{\max}^{2}+6\alpha^{2}R^{2}_{\max}+2(3\alpha^{2}+6\alpha^{4})\|Q^{*}\|^{2}$
$\hskip 28.45274pt+2(3\alpha^{2}+6\alpha^{4})\|\Psi_{n}\|^{2}$
$\leq K(1+\|\Psi_{n}\|^{2})\leq K(1+\|\Psi_{n}\|)^{2},$
[…]
4 Experiments
This section presents a comparison of the proposed two-step minimax Q-learning (TMQL) algorithm with the classical minimax Q-learning (MQL) [10], generalised optimal minimax Q-learning (G-SOROQL) [32], and generalised minimax Q-learning (G-SORQL) [32]. Throughout this section, TZMGs with $10$, $20$ and $50$ states with $5$ actions for each players are generated. The discount factor $\alpha$ is set to $0.6$. The algorithms are compared in terms of both error and time. Each episode of the experiment consists of running an algorithm for 1000 iterations. The error upon completion of the episode is calculated by taking the Euclidean norm difference between the optimal value and the estimated value upon completion of the episode. The average error is evaluated by taking the average over 50 independent episodes as follows
$\text{Average Error}=\dfrac{1}{50}\sum_{k=1}^{50}\|Y^{*}-val[Q_{k}(.)]\|_{2},$ (7)
where $Y^{*}$ is the min-max value function of the game and $val[Q_{k}(.)]$ is the min-max Q-value estimated upon completion of the $k^{th}$ episode using the particular algorithm. Similarly, the average time (in seconds) is calculated by taking the average over the time taken by the algorithm to complete an episode. The choice of parameters and problem set-up is similar to that of the experiment performed in [32]. It is essential to emphasize and acknowledge that the code utilized in this study has been sourced from the GitHub repository of R. B. Diddigi [53]. Our implementation can be accessed through the following link [54].
4-A MGs with the restriction on the structure
10 states 20 states 50 states
Algorithm
Average
Error
Average
Time
Average
Error
Average
Time
Average
Error
Average
Time
Standard minimax
Q-learning [10]
0.6819 1.571 1.670 1.458 3.9927 1.493
Generalised minimax
Q-learning [32]
0.5021 4.300 1.4550 6.303 3.7676 11.930
Generalised optimal minimax
Q-learning [32]
0.3364 1.592 1.2858 1.462 3.597 1.549
Two-step minimax
Q-learning
0.2184 2.201 1.170 2.122 3.532 2.232
TABLE I: Comparison of algorithms with varying states for $\alpha=0.6$, and $|A|=|B|=5$.
In this experiment, we first introduce a restriction on MGs. More specifically, we impose a condition on the transition probability of the generated MGs, such that $p(i|i,a,b)>0,\forall i,a,b$. This restriction is necessary for this experiment as the generalized minimax Q-learning algorithm (G-SORQL) [32], and generalized minimax optimal Q-learning algorithm (G-SOROQL) [32] developed only for MGs with this restriction. The choice of step-size sequence is same for all the experiments. The choice $\theta_{n}$ for this experiment is $\frac{80}{n+80}$.
Table I provides the comparison of all the algorithms for different states. In this table, each episode is run for 1000 iterations. Average error and average time is calculated over 50 episodes. In all the cases, the proposed method shows the superior performance in terms of average error. Note that G-SORQL algorithm involves evaluating the successive relaxation term at every iteration of the algorithm [32]. This increases the computational cost of the G-SORQL as the evaluation of successive relaxation term depends on the cardinality of the set $S$. This is evident in the time G-SORQL requires to finish an episode, as the cardinality of state $S$ increases. On the other hand the proposed TMQL, and existing MQL does not show much changes in the execution time as the state size increase. Therefore, one can observe the advantage of the proposed algorithm in terms of execution time in comparison to the G-SORQL.
In Figure 1, we plot the graph between number of iterations and average error for the choice of state space $S=10$, and action space $A=B=5$ with $\alpha=0.6$. The $x$-axis represents the number of iterations, while $y$-axis indicates the average error derived from $50$ independent episodes. Figure 1 depicts the convergence behaviour of all the algorithms. The figure clearly demonstrates that the average error is decreasing for all the algorithms. Moreover, it is evident that the proposed algorithm achieves faster convergence. Interestingly, even within this class of MGs, the proposed method outperforms the existing G-SORQL and G-SOROQL.
Figure 1: The algorithms performance over increasing iterations on examples of MGs with restrictions.
4-B Markov games (MGs) without restriction :
In this subsection, the choice of parameters is similar to the previous subsection, with the exception that we generate the examples without any restriction on the structure of the Markov games. More specifically, unlike [32] these examples with $10$ states, $20$ states, and $50$ states mentioned here has no restriction on $p(i|i,a,b)$. Also, the choice of the sequence $\theta_{n}$ for these examples is $\frac{100}{n+100}$.
Table II, provides the comparison of the proposed algorithm with standard minimax Q-learning [10] algorithm in terms of both error and time. It is evident from the table that the average error is significantly lower in comparison to MQL. Further, it is also evident that due to the extra computation involved in the TMQL, the time taken by the proposed algorithm is more compared to that of the classical MQL.
10 states 20 states 50 states
Algorithm
Average
Error
Average
Time
Average
Error
Average
Time
Average
Error
Average
Time
Standard minimax
Q-learning [10]
0.6843 1.1218 1.9466 1.141 3.954 1.198
Two-step minimax
Q-learning
0.3901 1.645 1.609 1.697 3.646 1.78
TABLE II: Comparison of MQL and TMQL, with varying states for $\alpha=0.6$ and $|A|=|B|=5.$
Remark 4.1.
As noted earlier in this example the proposed algorithm does not impose any restrictions on the structure of the MGs, distinguishing it from the approaches presented in [32]. In terms of practical utility, this characteristic is a significant improvement.
5 Conclusion
This manuscript presents a novel two-step RL algorithm for the two-player zero sum Markov game. The boundedness of the two-step minimax Q-learning and the convergence property of the algorithm is obtained theoretically. Effortlessly, one can implement the proposed two-step minimax Q-learning algorithm. Finally, empirical tests are conducted to confirm the advantage of the proposed algorithm.